# Specialization after Generalization: Towards Understanding Test-Time Training in Foundation Models

### v1

@@ C.3 Insufficiency of global training
We define a specific construction that satisfies Assumptions 1, 2, 3, 4 and 5 and models the superposition of concepts through a randomized feature map. This construction captures the challenge faced by an underparameterized model learning a complex environment.
@@
We first verify that Definition 7 satisfies Assumptions 1, 7, 8, 2, 3, 4 and 5, leading to a misspecification error of $\eta_{\text{mis}}=0$. As a consequence and by Proposition 3, TTT is consistent. Notably, the dimension of the feature map $d_{2}$ may be exponentially smaller than the number of concepts $d_{1}$, yet TTT can still learn the ground truth perfectly.
@@
    (Neighborhood preservation): We require $|\mathrm{sim}_{\Psi}(x,x^{\prime})-\mathrm{sim}_{\Phi}(x,x^{\prime})|\leq\eta_{\text{ang}}$. Consider ${x\in B_{i}},x^{\prime}\in B_{j}$ with $i\neq j$. $\mathrm{sim}_{\Phi}(x,x^{\prime})=\langle e_{i},e_{j}\rangle=0$. $\mathrm{sim}_{\Psi}(x,x^{\prime})=\langle p_{i},p_{j}\rangle$. Since $p_{i},p_{j}$ are drawn uniformly from $\mathcal{S}^{d_{2}-1}$, the inner product concentrates around 0. By standard concentration inequalities (related to the Johnson-Lindenstrauss lemma), with high probability over the draw of all pairs $\{p_{i}\}$, we have $\max_{i\neq j}|\langle p_{i},p_{j}\rangle|\leq O(\sqrt{\log(d_{1})/d_{2}})$. Thus, Assumption 2 holds with $\eta_{\text{ang}}$ small if $d_{2}$ is sufficiently large compared to $\log(d_{1})$.
@@
To verify Assumptions 4 and 5, we need to define the local linear maps $P_{m}$.
@@ C.4 Hardness of global learning under the LRH: main proposition and proof
Let $c_{1}\geq 0$, $\sigma\lesssim 1$, $k\asymp d^{c_{1}}$, and $s\asymp 1$. Assume $n\gtrsim\exp(\Omega(\log d)^{2})$. Then, Equation 5 holds, and there exists a polynomial-time algorithm $\mathcal{A}$ such that
@@
Under the low-degree polynomial conjecture, $n\gtrsim\exp(\Omega(\log d)^{2})$} samples are required by Algorithms based on Statistical Query (SQ) and the Low-Degree Polynomial (LDP) frameworks. Furthermore, without the low-degree polynomial conjecture, one can show that $\mathrm{poly}(n)$ samples are needed for any algorithm.
@@
Recall the definition of $\mathcal{F}_{d,s,k}$ and by using that $s\asymp 1$ and $\|f^{\star}\|_{L_{2}(\gamma_{d})}\asymp 1$, imply that
@@
$P_{t}[f^{\star}](z)=\mathbb{E}_{Z}[f^{\star}(\exp(-t)z+\sqrt{1-\exp(-2t)}Z]=\sum_{m=0}^{\infty}\exp(-tm)\mathcal{P}_{m}(f^{\star}),$
@@
Now, recall that we assume that $k$ is polynomial in $d$, and note that up to a measure zero, for $\Delta_{t}=P_{t}[f^{\star}]-f^{\star}$ it holds
@@
where we used our assumption that there are $k$ cells, the definition of the OU group, and the maximal inequality of $2k$-Gaussian. Therefore, we can choose $t\lesssim 1/\log(d)$, and obtain that
$\|P_{c/\log(d/n)}f^{\star}\|_{L_{2}(\gamma_{d})}\geq 1-c$
It implies that for $m\geq C\log(d)$, it holds that
$\|\mathcal{P}_{m}(f^{\star})\|_{L_{2}(\gamma_{d})}^{2}\lesssim\exp(-cm/\log(d))\|f^{\star}\|_{L_{2}(\gamma_{d})}^{2}\leq 0.1\cdot\exp(-cm/2),$
where we used that $P_{t+s}(f^{\star})=P_{t}P_{s}(f^{\star})$. In words, this property says that $f^{\star}$ is far from being a “low degree” polynomial. Meaning that,
@@
Therefore, we need to learn at most the top $\Theta(\log(d))$ basis of the Hermite polynomials. Using standard estimates from approximation theory (cf. Kalai et al. (2008)), samples to learn the $m=\Theta(\log(d))$-degree polynomials, let’s say by the projection estimator, which in words just means to learn the coefficient of the best low degree polynomial Bizeul & Klartag (2025).
@@
By definition, one can easily see that $\mathcal{F}_{d,s,k}$ is much easier to learn than the Gaussian Index Model (GIM). As the generative component of our functions are $k^{*}\asymp\log(d/n)$. It implies that under the low-polynomial degree conjecture, learning GIM requires $\exp(\Theta(\log(d/n)^{2}))$ samples, see Damian et al. (2024) and references within, and the seminal work of Arous et al. (2021). Without the low-degree polynomial conjecture, one may use the result of (Klivans et al., 2008, Thms. 26 and 27), which shows that the learnability of the subclass of
@@
requires at least $\exp(\Omega\log(d))\asymp d^{\Omega(1)}$ samples in the information theoretic sense. However, the algorithmic gap exists in those classical works Klivans et al. (2008); Kalai et al. (2008), as any efficient algorithm to learn this task requires $\exp(\Omega\log(d)^{2})$ samples.
@@ Appendix D Additional ablations
While the improvement of TTT in image classification may seem small in terms of classification error, we find that TTT can significantly reduce cross-entropy loss on the test set, as shown in Figure 7.
@@ Appendix E Experiment details
E.1 Comparison of logits (Figure 3)
@@ E.2 SAE on MNIST data
We use the same notation for active sets for concept vectors as in Section 3. However, for better clarity, we will recap the notation. For each sparse vector $\Phi(x)\in\mathbb{R}^{d_{1}}$ we define its active set as the set of vector components that are non-zero, i.e.,
@@
i.e., let mask $m$ only select indices that appear both in the test point and its neighbors? The flaw of this approach becomes evident for more complex data (e.g., ImageNet in Section 3), as slack components of the mask are necessary to select “non-spurious” test components.
@@ E.4 Scaling experiments on MNIST
In the scaling experiments reported in Section 4, we trained multiple LeNet (LeCun et al., 1998a) convolutional neural networks (CNNs) at different model scales. The architecture comprises two convolutional layers, each with ReLU activation and $2\times 2$ max-pooling, followed by a fully connected classification head. The reference models of varying sizes were trained with the hyperparameters listed in Table 3, obtained by tuning on the validation set. Optimization was performed with Adam (Kingma & Ba, 2015).
@@
For the model scaling plot in Figure 4, we trained each reference model across five random seeds and applied both TTT and majority voting. For TTT, all model parameters were frozen except for the final linear layer, which was fine-tuned from its pre-trained initialization using the hyperparameters in Table 4. Majority voting was performed with neighborhood sizes specified in Table 5. As the experiments were repeated over five seeds, we used 200 bootstrap iterations per seed to compute confidence intervals.
@@
To generate the dataset scaling plot in Figure 5, we randomly subsampled the training set while explicitly ensuring a uniform distribution across the 10 class labels. Here, TTT was performed by minimizing the cross-entropy loss over the $k=80$ nearest neighbors. Optimization was performed using Adam with a learning rate of 0.02 for 500 epochs.
We further conducted a mixture of experts (MoE) experiment, as shown in Figure 6, to empirically evaluate how well TTT can be approximated by a scalable MoE model. We used a pre-trained CNN to extract last-layer embeddings from MNIST images. Following Bertolissi et al. (2025), for a given number of experts, each expert was associated with a cluster centroid obtained by partitioning the training set into clusters via $k$-means. Expert training was performed by fine-tuning the pre-trained linear head on the nearest neighbors of the assigned centroid in embedding space. The fine-tuning hyperparameters are provided in Table 6. At test time, inputs were first mapped through the CNN encoder, after which the single closest cluster centroid was selected based on $L_{2}$-distance. The corresponding expert head then produced the final prediction.
@@
E.5 Scaling experiments on ImageNet
@@
Building upon the baseline linear classifier, we apply a test-time training (TTT) procedure to adapt the model at inference-time. Specifically, we fine-tune this linear head for each test sample by minimizing the cross-entropy loss over the set of $k=600$ nearest neighbors retrieved in the original CLIP embedding space. Optimization is carried out using the Adam optimizer with a learning rate of 0.02 and a batch size equal to the number of neighbors.
@@
To assess the impact of embedding dimensionality, we trained two-layer multi-layer perceptrons (MLPs) as classification heads on top of the CLIP embeddings. The first hidden layer had variable size, ranging from 250 to 4000 neurons, followed by a ReLU activation. The second layer maps the hidden representation to the 1000 ImageNet classes. Dropout is applied before the final layer to mitigate overfitting. Training was conducted with the Adam optimizer, using the hyperparameters specified in Table 7.
Hidden dimension
Hyperparameter 250 500 750 1000 1250 1500 2000 2500 3000 4000
Learning rate 4e-4 3e-4 3e-4 4e-4 3e-4 4e-4 4e-4 3e-4 4e-4 4e-4
@@
Table 7: Hyperparameters for globally training the MLP heads across different hidden dimensions.
After training, we freeze the first hidden layer and project the original 512-dimensional embeddings into this hidden space. The resulting hidden representations, taken after the ReLU activation, serve as our scaled embeddings for test-time training. Specifically, we fine-tune the pre-trained linear MLP head on the set of $k$ nearest neighbors by minimizing the cross-entropy loss using Adam. This setup allows us to systematically examine how performance and representation quality vary with embedding dimensionality. We used a set of universal hyperparameters that are nearly optimal across all hidden dimensions, as reported in Table 8. Loss optimization was performed in a full-batch setting, with the batch size equal to the number of neighbors.
@@
E.6 Scaling experiments for language modeling
We use the open-source implementation [^7] https://github.com/socialfoundations/tttlm of Hardt & Sun (2024) and evaluate the Qwen2.5 family of base models (Qwen et al., 2025). We summarize hyperparameters in Table 9. Sequences in the neighborhood that exceed the maximum sequence length are split into chunks of the maximum length. This means that a single neighbor can result in multiple gradient steps during TTT.

### v2

@@ C.3 Insufficiency of global training
We define a specific construction that satisfies Assumptions˜1, 2, 3, 4 and 5 and models the superposition of concepts through a randomized feature map. This construction captures the challenge faced by an underparameterized model learning a complex environment.
@@
We first verify that Definition 7 satisfies Assumptions˜1, 7, 8, 2, 3, 4 and 5, leading to a misspecification error of $\eta_{\text{mis}}=0$. As a consequence and by Proposition 3, TTT is consistent. Notably, the dimension of the feature map $d_{2}$ may be exponentially smaller than the number of concepts $d_{1}$, yet TTT can still learn the ground truth perfectly.
@@
    (Neighborhood preservation): We require $|\mathrm{sim}_{\Psi}(x,x^{\prime})-\mathrm{sim}_{\Phi}(x,x^{\prime})|\leq\eta_{\text{ang}}$. Consider ${x\in B_{i}},x^{\prime}\in B_{j}$ with $i\neq j$. $\mathrm{sim}_{\Phi}(x,x^{\prime})=\langle e_{i},e_{j}\rangle=0$. $\mathrm{sim}_{\Psi}(x,x^{\prime})=\langle p_{i},p_{j}\rangle$. Since $p_{i},p_{j}$ are drawn uniformly from $\mathcal{S}^{d_{2}-1}$, the inner product concentrates around 0. By standard concentration inequalities (related to the Johnson-Lindenstrauss lemma), with high probability over the draw of all pairs $\{p_{i}\}$, we have $\max_{i\neq j}|\langle p_{i},p_{j}\rangle|\leq O(\sqrt{\log(d_{1})/d_{2}})$. Thus, Assumption˜2 holds with $\eta_{\text{ang}}$ small if $d_{2}$ is sufficiently large compared to $\log(d_{1})$.
@@
To verify Assumptions˜4 and 5, we need to define the local linear maps $P_{m}$.
@@ C.4 Hardness of global learning under the LRH: main proposition and proof
Let $c_{1}\geq 0$, $\sigma\lesssim 1$, $k\asymp d^{c_{1}}$, and $s\asymp 1$. Assume $n\gtrsim\exp(\Omega(\log d)^{2})$. Then, Equation 5 holds, and there exists a polynomial-time algorithm $\mathcal{A}$ in its input, such that
@@
Under the low-degree polynomial conjecture, this bound is tight, $n\gtrsim\exp(\Omega(\log d)^{2})$ samples are required by any algorithm that relies on the Statistical Query (SQ) and the Low-Degree Polynomial (LDP) frameworks. Furthermore, without the low-degree polynomial conjecture, one can show that $\mathrm{poly}(n)$ samples are needed for any algorithm.
@@
Recalling the definition of $\mathcal{F}_{d,s,k}$ and using that $s\asymp 1$ and $\|f^{\star}\|_{L_{2}(\gamma_{d})}\asymp 1$, we obtain that
@@
$P_{t}[f^{\star}](x)=\mathbb{E}_{Z\sim\gamma_{n}}[f^{\star}(\exp(-t)x+\sqrt{1-\exp(-2t)}Z]=\sum_{m=0}^{\infty}\exp(-tm)\mathcal{P}_{m}(f^{\star}),$
@@
Now, recall that we assume $k$ is polynomial in $d$, and note that up to a measure zero, for $\Delta_{t}:=P_{t}[f^{\star}]-f^{\star}$, it holds for any $t\ll 1$ that
@@
where we used our assumption that there are $k$ cells, the definition of the OU group, and the maximal inequality of $2k$-Gaussian. Therefore, we can choose $t\leq c/\log(d)$, for small enough $c\geq 0$ and obtain that
$\|P_{c/\log(d/n)}f^{\star}\|_{L_{2}(\gamma_{d})}\geq 0.99.$
If there were more than $0.1$ $L_{2}^{2}(\gamma_{n})$, energy in the $m$-coefficients for $m\geq C\log(d/n)$ and $C\geq 0$ large enough, we would obtain a contraction. As for $t=c\log(d)$, it holds that
$\exp(-mt)\cdot\|\mathcal{P}_{m}(f^{\star})\|_{L_{2}(\gamma_{d})}^{2}\lesssim\exp(-cm/\log(d))\|f^{\star}\|_{L_{2}(\gamma_{d})}^{2}\leq 0.01\cdot\|\mathcal{P}_{m}(f^{\star})\|_{L_{2}(\gamma_{d})}^{2},$
which cannot align with the previous equation and our assumption of $\|f^{\star}\|_{L_{2}(\gamma_{n})}=1$. In words, this property says that $f^{\star}$ is far from being a “low degree” polynomial. Meaning that,
@@
Therefore, we need to learn at most the top $\Theta(\log(d))$ basis of the Hermite polynomials. Using standard estimates from approximation theory (cf. Kalai et al. (2008)), samples to learn the $m=\Theta(\log(d))$-degree polynomials, let’s say by the projection estimator, which in words means to learn the coefficients of the best low degree polynomial Bizeul & Klartag (2025).
@@
By definition, one can easily see that $\mathcal{F}_{d,s,k}$ is much harder to learn than the Gaussian Index Model (GIM). As the generative component of our functions are $k^{*}\asymp\log(d/n)$. It implies that under the low-polynomial degree conjecture, the aforementioned classes require at least $\exp(\Theta(\log(d/n)^{2}))$ samples to even learn GIM, see Damian et al. (2024) and references within, and the seminal work of Arous et al. (2021).
Without the low-degree polynomial conjecture, one may use the result of (Klivans et al., 2008, Thms. 26 and 27), which shows that the learnability of the subclass of
@@
requires at least $\exp(\Omega(\log(d))\asymp\mathrm{Poly}(d)$ samples in the information theoretic sense. However, the algorithmic gap exists in those classical works Klivans et al. (2008); Kalai et al. (2008), and a gap in the matching upper bound for the sample complexity.
@@ Appendix D Additional ablations
While the improvement of TTT in image classification may seem small in terms of classification error, we find that TTT can significantly reduce cross-entropy loss on the test set, as shown in Figure˜6.
@@ Appendix E Experiment details
E.1 Comparison of logits (Figure˜3)
@@ E.2 SAE on MNIST data
We use the same notation for active sets for concept vectors as in Section˜3. However, for better clarity, we will recap the notation. For each sparse vector $\Phi(x)\in\mathbb{R}^{d_{1}}$ we define its active set as the set of vector components that are non-zero, i.e.,
@@
i.e., let mask $m$ only select indices that appear both in the test point and its neighbors? The flaw of this approach becomes evident for more complex data (e.g., ImageNet in Section˜3), as slack components of the mask are necessary to select “non-spurious” test components.
@@ E.4 Scaling experiments on MNIST
In the scaling experiments reported in Section˜4, we trained multiple LeNet (LeCun et al., 1998a) convolutional neural networks (CNNs) at different model scales. The architecture comprises two convolutional layers, each with ReLU activation and $2\times 2$ max-pooling, followed by a fully connected classification head. The reference models of varying sizes were trained with the hyperparameters listed in Table˜3, obtained by tuning on the validation set. Optimization was performed with Adam (Kingma & Ba, 2015).
@@
For the model scaling plot in Figure˜4, we trained each reference model across five random seeds and applied both TTT and majority voting. For TTT, all model parameters were frozen except for the final linear layer, which was fine-tuned from its pre-trained initialization using the hyperparameters in Table˜4. Majority voting was performed with neighborhood sizes specified in Table˜5. As the experiments were repeated over five seeds, we used 200 bootstrap iterations per seed to compute confidence intervals.
@@
To generate the dataset scaling plot in Figure˜5, we randomly subsampled the training set while explicitly ensuring a uniform distribution across the 10 class labels. Here, TTT was performed by minimizing the cross-entropy loss over the $k=80$ nearest neighbors. Optimization was performed using Adam with a learning rate of 0.02 for 500 epochs.
E.5 Connection to MoEs
We used a pre-trained CNN to extract last-layer embeddings from MNIST images. Following Bertolissi et al. (2025), for a given number of experts, each expert was associated with a cluster centroid obtained by partitioning the training set into clusters via $k$-means. Expert training was performed by fine-tuning the pre-trained linear head on the nearest neighbors of the assigned centroid in embedding space. The fine-tuning hyperparameters are provided in Table˜6. At test time, inputs were first mapped through the CNN encoder, after which the single closest cluster centroid was selected based on $L_{2}$-distance. The corresponding expert head then produced the final prediction.
@@ E.5 Connection to MoEs
E.6 Scaling experiments on ImageNet
@@
Building upon the baseline linear classifier, we apply a test-time training (TTT) procedure to adapt the model at inference-time. Specifically, we fine-tune this linear head for each test sample by minimizing the cross-entropy loss over the set of $k=600$ nearest neighbors retrieved in the original CLIP embedding space. Optimization proceeds for 50 epochs using the Adam optimizer with a learning rate of 0.02 and a batch size equal to the number of neighbors.
@@
To assess the impact of embedding dimensionality, we trained two-layer multi-layer perceptrons (MLPs) as classification heads on top of the CLIP embeddings. The first hidden layer had variable size, ranging from 250 to 4000 neurons, followed by a ReLU activation. The second layer maps the hidden representation to the 1000 ImageNet classes. Dropout is applied before the final layer to mitigate overfitting. Training was conducted with the Adam optimizer, using the hyperparameters specified in Table˜7.
Model parameters
Hyperparameter 3.8e5 7.6e5 1.1e6 1.5e6 1.9e6 2.3e6 3.0e6 3.8e6 4.5e6 6.1e6
Hidden dimension 250 500 750 1000 1250 1500 2000 2500 3000 4000
Learning Rate 4.0e-4 3.5e-4 3.0e-4 4.0e-4 3.5e-4 4.0e-4 4.0e-4 3.5e-4 4.5e-4 4.5e-4
@@
Table 7: Hyperparameters for globally training the MLP heads across different model sizes.
After training, we freeze the first hidden layer and project the original 512-dimensional embeddings into this hidden space. The resulting hidden representations, taken after the ReLU activation, serve as our scaled embeddings for test-time training. Specifically, we fine-tune the pre-trained linear MLP head on the set of $k$ nearest neighbors by minimizing the cross-entropy loss using Adam. This setup allows us to systematically examine how performance and representation quality vary with embedding dimensionality. We used a set of universal hyperparameters that are nearly optimal across all model sizes, as reported in Table˜8. Loss optimization was performed in a full-batch setting, with the batch size equal to the number of neighbors.
@@
E.7 Scaling experiments for language modeling
We use the open-source implementation [^9] https://github.com/socialfoundations/tttlm of Hardt & Sun (2024) and evaluate the Qwen2.5 family of base models (Qwen et al., 2025). We summarize hyperparameters in Table˜9. Sequences in the neighborhood that exceed the maximum sequence length are split into chunks of the maximum length. This means that a single neighbor can result in multiple gradient steps during TTT.